OpenStudy (anonymous):

What is (3x+1)^2 simplified

OpenStudy (angela210793):

=9x^2+6x+1 (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

ask your math teacher why \[(3x+1)^2\] is more or less "simple" than \[9x^2+6x+1\]

OpenStudy (anonymous):

ok maybe don't. but i would love to know the answer. for example if i wanted to solve \[9x^2+6x+1=0\] it would surely be a lot "simpler" if i started with \[(3x+1)^2=0\] there is no such thing as "simpler" in mathematics, especially not in this case!

OpenStudy (angela210793):

hahha lol :D :D Yea I want to know the answer of tht q too :)